If the measures of two complementary angles are 16z-9 and 4z+3, What are the measures of the angles?
Answer by checkley71(8403)

16z-9+4z+3=90
20z=90-3+9
20z=96
z=96/20
z=4.8
therefore the 2 angles are:
16*4.8-9=67.8 &
4*4.8+3=22.2
proof
67.8+22.2=90
90=90